Here are simplified financial statements for Watervan Corporation: INCOME STATEMENT (Figures in $ millions) Net sales $ 891 Cost of goods sold 751 Depreciation 41 Earnings before interest and taxes (EBIT) $ 99 Interest expense 22 Income before tax $ 77 Taxes 26.95 Net income $ 50.05 BALANCE SHEET (Figures in $ millions) End of Year Start of Year Assets Current assets $ 379 $ 332 Long-term assets 278 232 Total assets $ 657 $ 564 Liabilities and shareholders’ equity Current liabilities $ 204 $ 167 Long-term debt 118 131 Shareholders’ equity 335 266 Total liabilities and shareholders’ equity $ 657 $ 564 The company’s cost of capital is 8%. a. Calculate Watervan’s economic value added (EVA). (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.) Economic value added $ million b. What is the company’s return on capital? (Use start-of-year rather than average capital.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) Return on capital % c. What is its return on equity? (Use start-of-year rather than average equity.) (Enter your answer as a percent rounded to 2 decimal places.) Return on equity % d. Is the company creating value for its shareholders? Yes No
Answer #1

Answer a:

Economic value added (EVA) = NOPAT - Capital charge

NOPAT = EBIT * (1 - Tax rate)

Tax rate = Taxes / Income before tax = 26.95 / 77 = 35%

NOPAT = 99 * (1 - 35%) = 64.35

Capital charge = Capital employed * Cost of capital = (Equity + Interest bearing liabilities) * Cost of capital

= (335 + 118) * 8%

= 36.24

Economic value added (EVA) = 64.35 - 36.24 = $28.11 million

Economic value added = $28.11 million

Answer b:

Company’s return on capital (using start-of-year capital) = NOPAT / (Total Assets - Current Liabilities)

= 64.35 /(564 - 167)

=16.21%

Return on capital = 16.21%

Answer c:

Return on equity (Using start-of-year equity) = Net Income / Shareholders' equity = 50.05 / 266 = 18.82%

Return on equity = 18.82%

Answer d:

Yes.

The company is creating value for its shareholders. The EVA for the year is $28.11 million. Company's Return on equity and Return on capital (both) are higher than cost of capital.